Circular No. 6207

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 41P/TUTTLE-GIACOBINI-KRESÁK S. Nakano, Sumoto, Japan, has forwarded observations by H. Abe

(Yatsuka, Japan, 0.26-m reflector + CCD), showing that this comet is in outburst, at m_{1} = 9.0 on Aug. 18.458 UT. These observations, made at the request of the Central Bureau, confirmed that the sighting on Aug. 17.04 by L. Shotter, Uniontown, PA, of a possible bright comet close to the predicted location of 41P was indeed that comet. Shotter estimated m_{1} ∼ 8 (0.15-m refractor; coma diameter ∼ 3′.5 with slight condensation). This comet also experienced two large outbursts in 1973 (cf. IAUC 2541, et seq.).

COMET 1995 Q1 The following position was measured by R. M. West, European South-

ern Observatory, from the exposure taken by G. Pizarro (cf. IAUC 6206):

1995 UT  α_{2000}  δ_{2000}
Aug. 17.99097  11^{h}02^{m}20.^{s}9  −13^{o}28′45′′

GRO J1735–27 R. B. Wilson, S. N. Zhang, M. Scott, and B. A. Harmon, Marshall

Space Flight Center, NASA; and T. Koh, D. Chakrabarty, B. Vaughan, and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE team: “Pulsations from a previously-uncatalogued source were detected on July 29 in the galactic-center region by BATSE. The observed period history shows a clear orbital signature. The maximum pulsed intensity was observed on Aug. 13, with a measured barycentric pulse period of 4.45565 ± 0.000025 s; the pulses continue to be observed as of Aug. 16, at energies from 20 to at least 70 keV. Using occultation imaging, we observed a source of ∼ 100 mCrab in the energy range 20–50 keV not present during July 2–14. A preliminary location (90-percent confidence) is bounded by a quadrilateral with vertices (equinox 2000.0) at α = 17^{h}36^{m}, δ = −26^{o}.9; 17^{h}37^{m}, −28^{o}.0; 17^{h}33^{m}, −27^{o}.3; 17^{h}32^{m}, −26^{o}.1. Positions estimated independently from the pulsed and DC flux are consistent at a 68-percent confidence level. If the detected emissions are from the same source, the pulsed fraction is ∼ 30 percent. The energy spectrum of the pulsed component on Aug. 5 was well fitted by a power-law model F (E) = C [E/(30 keV)]^{−α}, with C = (2.7 ± 1.0) × 10^{−4} photon cm^{−2} s^{−1} keV^{−1}, and α = 3.5 ± 1.0.”

1995 August 18 Daniel W. E. Green